a screw jack with pitch 0.5cm and the handle of the screw jack is 2cm long.Determine the force(effort)that must be applied at the end when lifting a load of 200N.Consider the efficiency of the screw to be 35%.


1
Expert's answer
2021-06-03T09:15:33-0400

One complete rotation of the handle of the screw jack leads to 0.5 cm vertical displacement of the load. So, the velocity ratio is


"VR=\frac{2\pi r}{p}."

The mechanical advantage is


"MA=\eta VR=\frac{2\pi\eta r}{p}."

So, the applied force is


"F_a=\frac{F}{MA}=\frac{Fp}{2\pi\eta r},\\\space\\
F_a=\frac{200·0.005}{2\pi·0.35·0.02}=22.7\text{ N}."